PCR Question: after 25 cycles in the thermometer, the DNA template is amplified to _____ copies.

Choices

a) 25 to the power of 2
b) over 100,000
c) over a million
d) over 33 million
Answer #1

Copies of DNA after n number of cycles after PCR = 2^n

Here, n = 25

Copies of DNA = 2^25 = 33,554,432

Option D

Please rate high.